OpenStudy (anonymous):

What is the area of a circle with circumference of 37.68? (Use 3.14 for p.)

OpenStudy (anonymous):

Circumference = 2*pi*r 37.68 = 2*3.14*r r = (37.68)/(6.28) r = 6 Area = pi*r^2 = 3.14*6*6 =113.04